OpenStudy (happyfeet8040):

WILL GIVE MEDALLL The table below shows the distance d(t) in feet that an object travels in t seconds: What is the average rate of change of d(t) between 3 seconds, and 5 seconds and what does it represent? 59.5 ft/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 ft/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 ft/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds 59.5 ft/s; it represents the average distance traveled by the object betwe

OpenStudy (happyfeet8040):

Please helppp

OpenStudy (anonymous):

Before I begin to answer your question, what approaches have you made to solve this problem?

OpenStudy (happyfeet8040):

Well, I was thinking it was C because it's asking how much distance has the object traveled within 3-5 seconds.. Am I right?

OpenStudy (anonymous):

No, since rate is units/ unit of time, that means our final result in this problem will be in ft/s. What property is in units of distance per unit time?

OpenStudy (happyfeet8040):

Okay, so B? It would be in speed instead of distance?

OpenStudy (anonymous):

Yes. The reason is \[speed = \frac {distance}{time}\] Therefore, since we want to know the average rate of change from second 3 to second five, our formula is \[speed = \frac {350 ft - 126 ft}{5s - 3 s} = 112 \frac {ft}{s}\]

OpenStudy (anonymous):

Remember, average rate of change simply means \[rate = \frac {unit}{time}\]

OpenStudy (happyfeet8040):

Oh, okay! That makes sense, I understand now. Thank you so much :)

OpenStudy (anonymous):

No problem

OpenStudy (happyfeet8040):

Can you help me with a couple more maybe?